Sol Summary – November 16th

Sol: 3
Summary Title: Intertribal Space conference
Author’s name: Izabela Shopova
Mission Status: Nominal
Sol Activity Summary:
We had a late start today giving us an extra hour of sleep in the morning to compensate for the late night, watching the recorded Artemis launch. After breakfast, we completed the first part of the VR First Aid training and our commander supported the Intertribal conference as planned. There were no questions from the conference. The afternoon EVA highlighted some communication issues which we discussed in the debriefing before dinner. Crew scientists collected samples for research and the crew took amazing photos.
Look Ahead Plan: Two EVAs tomorrow as planned and continued work on other projects
Anomalies in work: Radio contact was lost during EVA. The crew on EVA was able to communicate with each other but CapCom couldn’t hear them.
Weather: Cold, sunny and clear.
Crew Physical Status: Good. The crew is making an effort to drink more water and keep hydrated.
EVA: Successful afternoon EVA to White Rock canyon for biological samples collection.
Reports to be filed:
Sol Summary
Operations Report
Commander’s Report
EVA Request
EVA Report
Support Requested: None

Thank you Mission support

Operations Report – November 16th

SOL: 3
Name of person filing report: Jas Purewal
Non-nominal systems: Pipe under sink leaked a bit because it isn’t attached properly. There is no seal. For now we have left it and are going to drain the sink very slowly.
Notes on non-nominal systems: NA
Spirit rover used: Yes
Hours: (before EVA) 197.7
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging) 62%
Currently charging: No
Opportunity rover used: No
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: No
Curiosity rover used: Yes
Hours: 209.9
Beginning charge: 100%
Ending charge: 67%
Currently charging: No
Perseverance rover used: NA
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: No
General notes on rovers: All working fine
Summary of Hab operations: put notes here
WATER USE: 15 gallons
Water (static tank): 410 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off) off
Toilet tank emptied: yes
Summary of internet: Good
Summary of suits and radios:
Summary of GreenHab operations: None
WATER USE: 5.02 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: None
Dual split: (Heat or AC, On or Off) Heat, On
Summary of RAM operations: (Tools used, work done) Two GPS units taken to main hab for use. Didn’t end up using them on today’s EVA. But we may use them tomorrow, if we don’t we will return them to the RAM. Replaced 3AAA batteries on one of the GPS units.
Summary of any observatory issues: NA
Summary of health and safety issues: Nose bleeds have improved/reduced.
Questions, concerns and requests to Mission Support: Still waiting on more trash bags please for toilet bins and for general bins. Pillows please.

GreenHab Report – November 16th

Sol 3
GreenHab Officer: Caitlyn Hubric
Environmental control: (heater, fan, door open?) heater on, door closed, 20% humidity
Average temperatures: 43.3ºC/13.5ºC
Hours of supplemental light: N/A
Daily water usage for crops: 18 L ≈ 4.76 gal
Daily water usage for research and/or other purposes: 1 L ≈ 0.26 gal
Water in Blue Tank ~288 gallons
Time(s) of watering for crops: twice, morning and night, 9am-6pm
Changes to crops: none
-Tomorrow morning I will harvest the radish microgreens, I will taste the unlabeled bin and try to guess what it is and then harvest them, and I will provide support to the beans
Harvest: (include which crop and mass in grams)
Support/supplies needed:

EVA Report – November 16th

EVA # 3
Author: Elizabeth Balga
Purpose of EVA: Prospecting for and initial biological sample collection
Start time: 1:30 PM
End time: 4:00 PM (actual end: 3:30 PM)
Narrative: Three crew members embarked on the first biological sample collection EVA. Upon approach of “Kissing Camel Ridge ”, the crew experienced difficulties with the communication system when communicating locally between EVA participants. After consulting with HABCOM and the OUTPOST, the crew switched to Channel 2 and local communications issues were resolved. With concurrence from HABCOM, the crew proceeded to their destination White Rock Canyon. Along the road, Martian life (cows) and beautiful views were enjoyed. At the destination, BIO trained XO and HSO on proper sample collection and labeling practices. The crew collected eleven organic and inorganic samples including live plants, and BIO scouted several additional sites for future sample collection.
Destination: White Rock Canyon
Coordinates (use UTM WGS 84): 520000, 4248000
Participants: XO, BIO, HSO
Road(s) and routes per MDRS Map: Cow Dung Rd (0110) to White Rock Canyon on Rover
Mode of travel: Spirit & Curiosity

Commander Report – November 16th

Sol: 3
Summary Title: Break a Leg
Author’s name: Dr. Jenni Hesterman, Commander
Mission Status: Nominal
Commander Report: The crew opted to sleep in an hour this morning after viewing a delayed video and convened for breakfast at 8:30AM for strawberry pancakes and yogurt.

First up on the agenda was first aid training using an Oculus VR headset. After viewing the training, the crew practiced splinting a broken leg using material found in the habitat. We’re learning that everything in the station must have multiple uses since resources are limited and resupply is unknown.

The crew then supported the Intertribal Space Conference with a video explaining our mission and stood ready to address questions with the appropriate communication delay between Earth and Mars.

After finishing dinner leftovers for lunch, three crew members left on two rovers for a 2 hour EVA to explore White Rock Canyon. The crew initially experienced a communications problem in an area called Kissing Camel Ridge and learned they needed to move to a different radio channel. After sorting out this issue, they proceeded to the area and searched for unique plants, lichens and rocks. The crew collected eleven organic and inorganic samples and scouted several additional sites for future sample collection.

When EVA operations were complete, the crew met to discuss lessons learned from the mission. After a dinner of chicken and biscuits, we finalized the next day’s schedule. Priorities tonight include preparing for two EVAs tomorrow and a group learning preference and communication exercise.

Ad astra!

Copyright © The Mars Society. All rights reserved. | Main Site